OpenStudy (anonymous):

A sequence is defined by the recursive formula f (n + 1) = f(n) – 2. If f(1) = 18, what is f(5)?

OpenStudy (xapproachesinfinity):

f(2)=f(1)-2=18-2=16 f(3)=f(2)-2=16-2=14 f(4)=f(3)-2=14-2=12 and so on